OpenStudy (anonymous):

I really really need help with this question and its timed. I cant figure it out! Please help. A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

Please , i really need help. Its timed and i am absolutely clueless on how to do this. I am reading over my textbook over and over again and i cant figure it out.

OpenStudy (anonymous):

Okay' I'll try solving it if I can.

OpenStudy (anonymous):

Okay, thanks so much

OpenStudy (anonymous):



OpenStudy (anonymous):

Now try this. 1. Sin30 = BC/AB Recall than Sin30 is 1/2, substitute the value, 1/2 = 41/AB Cross multiply AB = 2*41 = 82 feet So, AB = 82 feet. So the length of the hypotneuse is 82 feet. 2. Take Sin60 = AC/AB Recall that Sin60 = Square root 3/2 Substitute the value down Square root 3/2 = AC/82 Cross multiply 82*Square Root 3/2 = AC 41*Square Root 3 = AC Know that Square Root of 3 = 1.73 Substitute 41*1.73 = 70.93 = AC So, AC = 70.93 feet

OpenStudy (anonymous):

a. So, length of the side opposite to the 60 degree angle = 70.93 feet. b. Length of the hypotenuse of the triangular = 82 feet.

OpenStudy (anonymous):

c. Sin30 = 41/82 = 1/2 = 0.5 Cos30 = 70.93/82 = 0.8650 Tan30 = 41/70.93 = 0.5780

OpenStudy (anonymous):

This is the best way I think I can solve it.

OpenStudy (anonymous):

Okay, thanks . I just dont understand the whole conspet of it

OpenStudy (anonymous):

Since it's the special 30 - 60 right triangle Hypo. length : 2 * 41 = 82 ft The longer leg: √3 * 41 = 70 ft

OpenStudy (anonymous):

Comparing with: sin 30 = Op./ hyp = 1/2 -> hyp = 2 * 41 = 82 ft Cos30 = Adj / hyp = √3/2 -> Adj = √3/2 * 82 = 70 ft Tan30 = Op/ Adj = √3/3